# Identifying Weight-Variant Latent Causal Models

### v6

@@ Abstract
The task of causal representation learning aims to uncover latent higher-level causal representations that affect lower-level observations. Identifying true latent causal representations from observed data, while allowing instantaneous causal relations among latent variables, remains a challenge, however. To this end, we start with the analysis of three intrinsic indeterminacies in identifying latent space from observations: transitivity, permutation indeterminacy, and scaling indeterminacy. We find that transitivity acts as a key role in impeding the identifiability of latent causal representations. To address the unidentifiable issue due to transitivity, we introduce a novel identifiability condition where the underlying latent causal model satisfies a linear-Gaussian model, in which the causal coefficients and the distribution of Gaussian noise are modulated by an additional observed variable. Under some mild assumptions, we can show that the latent causal representations can be identified up to trivial permutation and scaling. Furthermore, based on this theoretical result, we propose a novel method, termed Structural caUsAl Variational autoEncoder (SuaVE), which directly learns latent causal representations and causal relationships among them, together with the mapping from the latent causal variables to the observed ones. Experimental results on synthetic and real data demonstrate the identifiability and consistency results and the efficacy of SuaVE in learning latent causal representations.
Code is available at : https://sites.google.com/view/yuhangliu/projects
Keywords: Causal Representations, Latent Causal Models, Identifiability
@@ 1 Introduction
While there is no universal formal definition, one widely accepted feature of disentangled representations (Bengio et al., 2013) is that a change in one dimension corresponds to a change in one factor of variation in the underlying model of the data, while having little effect on others. The underlying model is rarely available for interrogation, however, which makes learning disentangled representations challenging. Several excellent works for disentangled representation have been proposed that focus on enforcing independence over the latent variables that control the factors of variation (Higgins et al., 2017; Chen et al., 2018; Locatello et al., 2019; Kim and Mnih, 2018; Locatello et al., 2020). In many applications, however, the latent variables are not statistically independent, which is at odds with the notion of disentanglement, i.e., foot length and body height exhibit strong correlation in the observed data (Träuble et al., 2021).
@@
Causal representation learning avoids the aforementioned limitation, as it aims to learn a representation that exposes the unknown high-level causal structural variables, and the relationships between them, from a set of low-level observations (Schölkopf et al., 2021). Unlike disentangled representation learning, it identifies the possible causal relations among latent variables. In fact, disentangled representation learning can be viewed as a special case of causal representation learning where the latent variables have no causal influences (Schölkopf et al., 2021). One of the most prominent additional capabilities of causal representations is the ability to represent interventions and to make predictions regarding such interventions (Pearl, 2000), which enables the generation of new samples that do not lie within the distribution of the observed data. This can be particularly useful to improve the generalization of the resulting model. Causal representations also enable answering counterfactual questions, e.g., would a given patient have suffered heart failure if they had started exercising a year earlier?
@@
Despite its advantages, causal representation learning is a notoriously hard problem—without certain assumptions, identifying the true latent causal model from observed data is generally not possible. There are four primary approaches to achieve identifiability: 1) adapting (weakly) supervised methods with given latent causal graphs or/and labels (Kocaoglu et al., 2018; Yang et al., 2021; Von Kügelgen et al., 2021; Brehmer et al., 2022), 2) imposing special graphical conditions, e.g., with bottleneck graphical conditions (Adams et al., 2021; Xie et al., 2020a; Lachapelle et al., 2021), 3) using temporal information (Yao et al., 2021; Lippe et al., 2022), 4) using hard interventional data (Ahuja et al., 2023; Seigal et al., 2022). A brief review is provided in Section 2. For the supervised approach, when labels are known, the challenging identifiability problem in latent space has been transferred to an identifiability problem in the observed space, for which some commonly-used functional classes have been proven to be identifiable (Zhang and Hyvarinen, 2012; Peters et al., 2014). Given latent causal graphs overly depends on domain knowledge. For the second approach, many true latent causal graphs do not satisfy the assumed special graph structure. The temporal approach is only applicable when temporal information or temporal intervened information among latent factors is available. For using intervention data, most studies focus on hard interventions, particularly single-node and paired hard interventions. In contrast, our work considers the change of causal influences, which is more aligned with soft interventions.
In this work, we explore a new direction in the identifiability of latent causal representations by examining the change in causal influences among latent causal variables, motivated by recent advances in nonlinear ICA (Hyvarinen et al., 2019; Khemakhem et al., 2020). Hyvarinen et al. (2019); Khemakhem et al. (2020) have shown that with an additional observed variable $\mathbf{u}$ to modulate latent independent variables, the latent independent variables conditional on $\mathbf{u}$ are identifiable. This raises a natural question: beyond the independent latent variables in nonlinear ICA, with causal relationships among latent variables, what additional assumptions are required for identifiability of latent causal variables? To answer the question above, we start from the analysis of three intrinsic indeterminacies in latent space (see Section 3): transitivity, permutation indeterminacy, and scaling indeterminacy, which further give rise to the following insights. 1) Transitivity is the scourge of identifiability of latent causal models. 2) Permutation indeterminacy nature enables us to pre-define a causal order, avoiding troublesome directed acyclic graph (DAG) constraints. 3) Scaling and permutation indeterminacy only allow recovering latent causal variables up to permutation and scaling, not the exact values. To overcome the challenge of transitivity, we model the underlying causal representation with weight-variate linear Gaussian models (see Section 4). In this model, both the weights (i.e., causal influences) and the mean and variance of the Gaussian noise are modulated by an additional observed variable $\mathbf{u}$. By exploring the change in causal influences, we demonstrate that the latent causal representations can be recovered up to a trivial permutation and scaling. Combined with the independent causal mechanism, a common assumption in causality, we show that the latent causal structure is also identifiable. Further, realizing that requiring all weights to change might be challenging in real applications, we show partial identifiability results if only part of the weights change in Section 5. Based on our analysis, in Section 6, we additionaly propose a novel method, Structural caUsAl Variational autoEncoder (SuaVE), for learning latent causal representations and latent causal graph among causal representations. Section 7 verifies the efficacy of the proposed approach on both synthetic and real fMRI data.
@@
The key to the identifiability result in this work lies in exploring the changes in causal influences (weights) among the latent causal variables. Essentially, these changes correspond to distribution shifts, which can be viewed as natural interventions acting on latent causal variables in the context of causal representation learning. Such shifts frequently occur in various fields, including medical imaging (Chandrasekaran et al., 2021), biogeography (Pinsky et al., 2020), and finance (Gibbs and Candes, 2021). Similar to the exploration of distribution shifts in observed space (Ghassami et al., 2018; Huang et al., 2020), distribution shifts in latent space also provide opportunities for comparative analysis, offering insights into the underlying causal mechanisms governing the relationships between latent variables. Therefore, investigating distribution shifts for the identifiability of causal representations is a promising approach.
@@ 2 Related work
Due to the challenges of identifiability in causal representation learning, most existing works handle this problem by imposing assumptions. We thus give a brief review of the related work on this basis.
@@
Approaches falling within this class assume known latent causal graphs or labels. CausalGAN (Kocaoglu et al., 2018) requires a-priori knowledge of the structure of the causal graph of latent variables, which is a significant practical limitation. CausalVAE (Yang et al., 2021) needs additional labels to supervise the learning of latent variables. Such labels are not commonly available, however, and manual labeling can be costly and error-prone. Von Kügelgen et al. (2021) use a known but non-trivial causal graph between content and style factors to study self-supervised causal representation learning. Brehmer et al. (2022) learn causal representation in a weakly supervised setting whereby they assume access to data pairs representing the system before and after a randomly chosen unknown intervention.
Sparse Graphical Structure
Most recent progress in identifiability focuses on sparse graphical structure constraints (Silva et al., 2006; Shimizu et al., 2009; Anandkumar et al., 2013; Frot et al., 2019; Cai et al., 2019; Xie et al., 2020a, 2022a). Adams et al. (2021) provided a unifying viewpoint of this approach whereby a sparser model that fits the observation is preferred. However, they primarily consider linear relations between latent causal variables, and between latent causal variables and observed variables, which are often violated in real-world applications. Lachapelle et al. (2021) handles nonlinear causal relation between latent variables by assuming special sparse graphical structures. However, many latent causal graphs in reality may be more or less arbitrary, beyond a purely sparse graph structure. In contrast, our work assumes a function class of latent variables, and does not restrict the graph structure over them.
@@
The temporal constraint that the effect cannot precede the cause has been used repeatedly in latent causal representation learning (Yao et al., 2021; Lippe et al., 2022; Yao et al., 2022). For example, Yao et al. (2021) recover latent causal variables and the relations between them using Variational AutoEncoders and enforcing constraints in causal process prior. Lippe et al. (2022) learn causal representations from temporal sequences, which requires the underlying causal factors to be intervened. All of these works can be regarded as special cases of exploring the change of causal influences among latent variables in time series data. The approach proposed is more general since the observed auxiliary variable $\mathbf{u}$ could represent time indices, domain indices, or almost any additional or side information.
Hard Intervention
Very recently, there have been some works exploring interventional data (Brehmer et al., 2022; Ahuja et al., 2023; Seigal et al., 2022; Buchholz et al., 2023; Varici et al., 2023). Most of them consider hard intervention, and more restricted single-node and paired hard interventions (Ahuja et al., 2023; Seigal et al., 2022; Buchholz et al., 2023; Varici et al., 2023, 2024; Buchholz et al., 2024; von Kügelgen et al., 2024), which can only capture some special changes of causal influences. By contrast, this work studies unpaired data, and employs soft interventions to model a broader range of possible changes. This is especially important considering that distribution shifts in real applications may be arbitrary. From this perspective, soft interventions could be easier to achieve for latent variables than hard interventions.
Besides, Kivva et al. (2021) considers a nonlinear setting for latent causal discovery. It assumes discrete latent causal variables that are rendered identifiable by a mixture oracle, while here we consider continuous latent causal variables.
@@ 3 Indeterminacies in Latent Causal Models
In this section, we first build a connection between nonlinear ICA and causal representation learning, by exploiting the correspondence between the independence of latent variables in nonlinear ICA and the independence of latent noise variables in causal representations. We then consider three indeterminacies in latent space: transitivity, permutation indeterminacy, and scaling indeterminacy, and analyze their impact on identifiability.
3.1 Relating Causal Representation with Identifiable Nonlinear ICA
Figure 1: Causal representation with $\mathbf{u}$.
Causal representation learning aims to uncover latent higher-level causal representations that can explain lower-level raw observations with a nonlinear mapping. Specifically, we assume that the observed variables $\mathbf{x}$ are influenced by latent causal variables ${z_{i}}$, and the causal structure among ${z_{i}}$ can be any directed acyclic graph (which is unknown). For each latent causal variable $z_{i}$, there is a corresponding latent noise variable $n_{i}$, as shown in Figure 1, which represents some unmeasured factors that influence $z_{i}$. The latent noise variables $n_{i}$ are assumed to be independent with each other, conditional on the observed variable $\mathbf{u}$ [^1] For convenience in the later parts, with a slight abuse of definition, independent $n_{i}$ means that $n_{i}$ are mutually independent conditional on the observed variable $\mathbf{u}$., in a causal system (Peters et al., 2017), so it is natural to leverage recent progress in nonlinear ICA (Hyvarinen et al., 2019; Khemakhem et al., 2020), which has shown that the independent latent noise variables $n_{i}$ are identifiable under relatively mild assumptions, i.e., one of the main assumptions is that $\{n_{i}\}$ are Gaussian distributed with their mean and variance modulated by an observed variable $\mathbf{u}$. Taking one step further, our goal is to recover the latent causal variables $z_{i}$. However, with the assumptions for the identifiability of latent noise variables $n_{i}$ from nonlinear ICA, it is still insufficient to identify the latent causal variables $z_{i}$. The reason of such non-identifiability will be given in Section 3.2. Given this fact, a further question is what additional conditions are needed to recover the latent causal variables $z_{i}$. The corresponding identifiability conditions will be given in Section 4.
@@ 3.1 Relating Causal Representation with Identifiable Nonlinear ICA
Essentially, the introduced $\mathbf{u}$ as a surrogate variable is to characterize the changes of independent noise variables $n_{i}$ conditional on $\mathbf{u}$. In some cases, such a surrogate variable may be unavailable or hard to obtain in practice Liu et al. (2021); Creager et al. (2021). Fortunately, it has been shown that learning or inferring such a surrogate variable is possible with additional auxiliary information Lin et al. (2022). These information is often cheaply available for every input in practice Xie et al. (2020b); Wang et al. (2020). Examples include time index of the data in time series forecasting tasks Mudelsee (2019), locations (longitude and latitude) of collected satellite data in remote sensing Rußwurm et al. (2020), modality index in multi-modality dataset, and label in natural images.

### v7

@@ Abstract
The task of causal representation learning aims to uncover latent higher-level causal variables that affect lower-level observations. Identifying the true latent causal variables from observed data, while allowing instantaneous causal relations among latent variables, remains a challenge, however. To this end, we start with the analysis of three intrinsic indeterminacies in identifying latent variables from observations: transitivity, permutation indeterminacy, and scaling indeterminacy. We find that transitivity acts as a key role in impeding the identifiability of latent causal variables. To address the unidentifiable issue due to transitivity, we introduce a novel identifiability condition where the underlying latent causal model satisfies a linear-Gaussian model, in which the causal coefficients and the distribution of Gaussian noise are modulated by an additional observed variable. Under certain assumptions, including the existence of a reference condition under which latent causal influences vanish, we can show that the latent causal variables can be identified up to trivial permutation and scaling, and that partial identifiability results can still be obtained when this reference condition is violated for a subset of latent variables. Furthermore, based on these theoretical results, we propose a novel method, termed Structural caUsAl Variational autoEncoder (SuaVE), which directly learns causal representations and causal relationships among them, together with the mapping from the latent causal variables to the observed ones. Experimental results on synthetic and real data demonstrate the identifiability and consistency results and the efficacy of SuaVE in learning causal representations.
Project Page:
Keywords: Causal Representation Learning, Latent Causal Models, Nonlinear ICA
@@ 1 Introduction
While there is no universal formal definition, one widely accepted feature of disentangled representations is that a change in one dimension corresponds to a change in one factor of variation in the underlying model of the data, while having little effect on others (Bengio et al., 2013). The underlying model is rarely available for interrogation, however, which makes learning disentangled representations challenging. Several excellent works for disentangled representation have been proposed that focus on enforcing independence over the latent variables that control the factors of variation (Higgins et al., 2017; Chen et al., 2018; Locatello et al., 2019; Kim and Mnih, 2018; Locatello et al., 2020). In many applications, however, the latent variables are not statistically independent, which is at odds with the notion of disentanglement, i.e., foot length and body height exhibit strong correlation in the observed data (Träuble et al., 2021).
@@
Causal representation learning avoids the aforementioned limitation, as it aims to learn a representation that exposes the unknown high-level latent causal variables, and the relationships between them, from a set of low-level observations (Schölkopf et al., 2021). Unlike disentangled representation learning, it allows the possible causal relations between latent variables. In fact, disentangled representation learning can be viewed as a special case of causal representation learning where the latent variables have no causal influences (Schölkopf et al., 2021). One of the most prominent additional capabilities of causal representations is the ability to represent interventions and to make predictions regarding such interventions (Pearl, 2000), which enables the generation of new samples that do not lie within the distribution of the observed data. This can be particularly useful to improve the generalization of the resulting model. Causal representations also enable answering counterfactual questions, e.g., would a given patient have suffered heart failure if they had started exercising a year earlier?
@@
Despite its advantages, causal representation learning is a notoriously hard problem. Without certain assumptions, identifying the true latent causal model from observed data is generally not possible. There are four primary approaches to achieve identifiability: 1) adapting (weakly) supervised methods with given latent causal graphs or/and labels (Kocaoglu et al., 2018; Yang et al., 2021), 2) imposing special graphical conditions with bottleneck graphical conditions (Adams et al., 2021; Xie et al., 2020a; Lachapelle et al., 2021), 3) using temporal information (Yao et al., 2021; Lippe et al., 2022), 4) using hard interventional data (Ahuja et al., 2023; Seigal et al., 2022). A brief review is provided in Section 2. In supervised approaches, when either the labels for latent variables or the true latent causal graph are available, the challenging identifiability problem in the latent space can be alleviated, or even bypassed by leveraging existing identifiability results in observed sapce (Zhang and Hyvarinen, 2012; Peters et al., 2014). However, such methods tend to rely heavily on domain-specific prior knowledge and are often associated with high annotation costs. For the second approach, many true latent causal graphs do not satisfy the assumed special graph structure. The temporal approach is only applicable when temporal information or temporal intervened information between latent factors is available. For using intervention data, most studies focus on hard interventions, particularly single-node and paired hard interventions. In contrast, our work considers the changes of causal influences, which is more aligned with soft interventions.
In this work, we explore a new direction in the identifiability of latent causal variables by examining the changes in causal influences between latent causal variables. Our approach is motivated by recent advances in nonlinear ICA, which, broadly speaking, demonstrate that latent independent variables can be identified when an additional observed variable $\mathbf{u}$ is introduced to modulate their distributions (Hyvarinen et al., 2019; Khemakhem et al., 2020). This raises a natural question:
Beyond the independent latent variables in nonlinear ICA, with causal relationships between latent variables, what additional assumptions are required for identifiability of latent causal variables?
  • To answer the question above, we begin by analyzing three intrinsic indeterminacies in the latent space (see Section 3): transitivity, permutation indeterminacy, and scaling indeterminacy. This analysis leads to the following key insights. 1) Transitivity is the scourge of identifiability of latent causal models. 2) Permutation indeterminacy nature enables us to pre-define a causal order, avoiding troublesome directed acyclic graph (DAG) constraints. 3) Scaling and permutation indeterminacy only allow recovering latent causal variables up to permutation and scaling, not the exact values.
  • To overcome the challenge of transitivity, we model the underlying causal generative process with weight-variate linear Gaussian models (see Section 4). In this model, both the weights (i.e., causal influences) and the mean and variance of the Gaussian noise are modulated by an additional observed variable $\mathbf{u}$.
  • By exploring the changes in causal influences, i.e., the existence of a reference environment under which certain causal influences between latent variables vanish, we demonstrate that the latent causal variables can be recovered up to a trivial permutation and scaling. Combined with the independent causal mechanism, a common assumption in causality, we show that the latent causal structure is also identifiable. Further, realizing that requiring all weights to change may be challenging in real applications, we show partial identifiability results if only part of the weights change in Section 5.
  • Based on our analysis, in Section 6, we additionally propose a novel method, Structural caUsAl Variational autoEncoder (SuaVE), for learning latent causal variables and latent causal graph among them. Section 7 verifies the efficacy of the proposed method on both synthetic and real fMRI data.
@@
The key to the identifiability results in this work lies in leveraging changes in the causal influences (e.g., weights in linear models) between latent causal variables. These changes correspond to distribution shifts, which can be interpreted as interventions acting on the latent variables. Such distribution shifts commonly occur across various domains, including medical imaging (Chandrasekaran et al., 2021), biogeography (Pinsky et al., 2020), and finance (Gibbs and Candes, 2021). Analogous to studies of distribution shifts in observed data (Ghassami et al., 2018; Huang et al., 2020), shifts in the latent space enable comparative analyses that reveal the underlying causal mechanisms governing relationships between latent variables. Thus, investigating distribution shifts offers a promising avenue for achieving identifiability in causal representation learning.
@@ 2 Related work
Due to the inherent challenges of identifiability in causal representation learning, most existing methods address this issue by imposing specific assumptions. We therefore provide a brief review of related work grounded on this basis.
@@
Approaches in this category assume either known latent causal graphs or access to labels. For example, CausalGAN (Kocaoglu et al., 2018) requires prior knowledge of the latent causal graph structure, which limits its practical applicability. CausalVAE (Yang et al., 2021) relies on additional labels to supervise latent variable learning. However, such labels are often unavailable, and manual annotation can be costly and error-prone. Von Kügelgen et al. (2021) explore self-supervised causal representation learning using a known but complex causal graph between content and style factors. Meanwhile, Brehmer et al. (2022) propose a weakly supervised approach that assumes access to paired data before and after unknown interventions on the system.
Special Graphical Structure
Most recent advances in identifiability focus on imposing special graphical structure constraints (Silva et al., 2006; Shimizu et al., 2009; Anandkumar et al., 2013; Frot et al., 2019; Cai et al., 2019; Xie et al., 2020a, 2022a; Kivva et al., 2021). Adams et al. (2021) provided a unifying perspective, advocating for sparser models that fit the observations better. However, their analysis mainly considers linear causal relationships both between latent causal variables and between latent and observed variables, which often violated in real-world complex scenarios. Lachapelle et al. (2021) addressed nonlinear causal relations by assuming special sparse graphical structures. Yet, latent causal graphs encountered in practice are more complex and may not conform to strict sparsity assumptions. In contrast, our approach defines a function class for latent variables without imposing restrictive constraints on the underlying graph structure. Additionally, Kivva et al. (2021) focus on discrete latent causal variables within relatively limited graph structures, while our work considers continuous latent causal variables with flexible general graph configurations.
@@
The temporal constraint that the effect cannot precede the cause has been used repeatedly in latent causal representation learning (Yao et al., 2021; Lippe et al., 2022; Yao et al., 2022). For example, Yao et al. (2021) recover latent causal variables and the relations between them using Variational AutoEncoders and enforcing constraints in causal process prior. Lippe et al. (2022) learn causal representations from temporal sequences, which requires the underlying causal factors to be intervened. All of these works can be regarded as special cases of exploring the change of causal influences among latent variables in time series data. The proposed approach in this work tends to be more general, since the auxiliary variable $\mathbf{u}$ can represent not only time indices but also domain indices or virtually any additional side information.
Data from Hard Interventions
Very recently, there have been some works exploring interventional data (Brehmer et al., 2022; Ahuja et al., 2023; Seigal et al., 2022; Buchholz et al., 2023; Varici et al., 2023). Most of them consider hard intervention, and more restricted single-node and paired hard interventions (Ahuja et al., 2023; Seigal et al., 2022; Buchholz et al., 2023; Varici et al., 2023, 2024; Buchholz et al., 2024; von Kügelgen et al., 2024), which can only capture some special changes of causal influences. By contrast, this work studies unpaired data, and employs soft interventions to model a broader range of possible changes. This is especially important considering that distribution shifts in real applications may be arbitrary. From this perspective, soft interventions could be easier to achieve for latent variables than hard interventions. To clarify this distinction, we introduce hard and soft interventions in the following.
Preliminaries: Hard Intervention and Soft Intervention
Generally speaking, hard intervention (sometimes called perfect intervention) fixes a causal variable to a specific value, thereby breaking its dependence on all of its parent variables. In contrast, soft intervention modifies the distribution of a causal variable conditional on its parent nodes, such as by altering its parameters or noise distribution, without completely severing the causal influence from its parents. This added flexibility allows soft interventions to capture a broader and more realistic range of changes in causal influences, enabling the model to account for a broader array of changes in causal influences. This is particularly valuable for capturing changes in causal influences driven by uncontrollable behaviors, such as environmental fluctuations, within a causal system. Moreover, soft interventions generalize hard interventions, since a hard intervention can be viewed as a special case of a soft intervention where the distribution of the targeted causal variable collapses to a Dirac delta distribution.
@@ 3 Indeterminacies in Latent Causal Models
In this section, we first build a connection between nonlinear ICA and causal representation learning, by exploiting the correspondence between the independence of latent variables in nonlinear ICA and the independence of latent noise variables in causal systems. We then consider three indeterminacies in latent space: transitivity, permutation indeterminacy, and scaling indeterminacy, and analyze their impact on identifiability.
3.1 Relating Causal Representation Learning with Identifiable Nonlinear ICA
Figure 1: Latent Causal Models with $\mathbf{u}$.
Causal representation learning aims to uncover latent higher-level causal variable that can explain lower-level raw observations. Specifically, we assume that the observed variables $\mathbf{x}$ are influenced by latent causal variables $z_{i}$}, and the causal structure among $z_{i}$} can be any directed acyclic graph (which is unknown). For each latent causal variable $z_{i}$, there is a corresponding latent noise variable $n_{i}$, as shown in Figure 1, which represents some unmeasured factors that influence $z_{i}$. The latent noise variables $n_{i}$ are assumed to be independent with each other, conditional on the observed variable $\mathbf{u}$ [^1] For convenience in the later parts, with a slight abuse of definition, independent $n_{i}$ means that $n_{i}$ are mutually independent conditional on the observed variable $\mathbf{u}$., in a causal system (Peters et al., 2017), so it is natural to leverage recent progress in nonlinear ICA (Hyvarinen et al., 2019; Khemakhem et al., 2020), which has shown that the independent latent noise variables $n_{i}$ are identifiable under relatively mild assumptions, i.e., one of the main assumptions is that $\{n_{i}\}$ are Gaussian distributed with their mean and variance modulated by an observed variable $\mathbf{u}$. Taking one step further, our goal is to recover the latent causal variables $z_{i}$. However, with the assumptions for the identifiability of latent noise variables $n_{i}$ from nonlinear ICA, it is still insufficient to identify the latent causal variables $z_{i}$. The reason of such non-identifiability will be given in Section 3.2. Given this fact, a further question is what additional conditions are needed to recover the latent causal variables $z_{i}$. The corresponding identifiability conditions will be given in Section 4.
@@ 3.1 Relating Causal Representation Learning with Identifiable Nonlinear ICA
Essentially, the introduced $\mathbf{u}$ as a surrogate variable is to characterize the changes of independent noise variables $n_{i}$ conditional on $\mathbf{u}$. In some cases, such a surrogate variable may be unavailable or hard to obtain in practice (Liu et al., 2021; Creager et al., 2021). Fortunately, it has been shown that learning or inferring such a surrogate variable is possible with additional auxiliary information (Lin et al., 2022). These information is often cheaply available for every input in practice (Xie et al., 2020b; Wang et al., 2020). Examples include time index of the data in time series forecasting tasks (Mudelsee, 2019), locations (longitude and latitude) of collected satellite data in remote sensing (Rußwurm et al., 2020), modality index in multi-modality dataset, and label in natural images.